OpenStudy (anonymous):

one leg of a right triangle is 5 times the other leg. find the longer leg if the hypotenuse is 60 ft. i just want the answer

OpenStudy (anonymous):

x^2 + (5x)^2 = 60^2 x^2 + 25x^2 = 360 26x^2 = 3600 x^2 = 138.46 x = 11.77 so the longer leg is 5x = 58.83